OpenStudy (anonymous):

Jersey-Cash-5 is a New Jersey Lottery that \bene ts education and institutions". In this lottery, a game is played by picking 5 distinct numbers out of numbers 1 to 40. Compute the number of ways and the probability of winning each of the following cases. Show the combinatorial derivation. Use a calculator to perform the nal step in the computation. Express the nal result as a ratio 1=x where the denomintor x is rounded to the nearest integer. (a) Jackpot: Match 5 out of 5. (b) Second Prize: Match 4 out of 5. (You win $500.) (c) Third Prize: Match 3 out of 5. (You win $11.)

OpenStudy (anonymous):

Please ans this question. exam in couple hours n this is likey to be on the exam !!

OpenStudy (anonymous):

@amistre64 and @TuringTest any idea abt this question ?

OpenStudy (amistre64):

what is the prob of picking any one number correctly?

OpenStudy (anonymous):

@amistre64 how do u define correctly ? i dint understand ur question

OpenStudy (amistre64):

if this is a fair game, there are 40 distinct elements to choose at random; so each element must have the same probability of being picked ... any given element therefore has a 1/40 probability of being chosen right?

OpenStudy (anonymous):

mm yes

OpenStudy (amistre64):

there is a success rate "s" of picking a correct number (1/40) this means that there is also a failure rate "f" of (39/40) there is a setup of choosing "n" numbers from these odds as: \[(s+f)^n\]

OpenStudy (amistre64):

this has been worked out as equivalent to saying:\[\Large\binom{n}{k}s^{k}f^{n-k}\]so in this case \[\Large\binom{5}{k}~\frac{1}{40}^{k}~\frac{39}{40}^{5-k}\] should give you the probability of choosing n numbers correctly out of 5

OpenStudy (amistre64):

let k be the number of correct choices that is

OpenStudy (anonymous):

i am not familiar with those formulas. can u tell me what are they and when are they used ?

OpenStudy (amistre64):

they are used for a discrete random variable that only has 2 likely outcomes - its called a binomial

OpenStudy (amistre64):

i cant see a simpler way to present it without having to write up a chapters worth of text :/

OpenStudy (anonymous):

hehe i understand ! is there any other way to do this problem without using this formulas?

OpenStudy (amistre64):

there are a certain number of ways to pick 5 elements; each draw is either a success or a failure 5C0 or 5C5 fffff sssss 5C1 or 5C4 ffffs ssssf fffsf sssfs ffsff ssfss fsfff sfsss sffff fssss 5C2 or 5C3 fffss sssff ffssf ssffs fssff sffss ssfff ffsss ffsfs ssfsf fsfsf sfsfs sfsff fsfss fsffs sfssf sffsf fssfs sfffs fsssf 32 actual outcomes

OpenStudy (amistre64):

the probability of getting any one of these is the product of their s and f values: s = 1/40 f = 39/10 this pretty much sums up the formula

OpenStudy (amistre64):

39/40 that is

OpenStudy (amistre64):

notice that there are 5C0 ways of getting none correct, s^0 and f^5 (5 0) (1/40)^0 (39/40)^5 notice that there are 5C5 ways of getting all correct, s^5 and f^0 (5 5) (1/40)^5 (39/40)^0 notice that there are 5C2 ways of getting 2 correct, s^2 and f^5-2 (5 2) (1/40)^2 (39/40)^3 etc ....

OpenStudy (anonymous):

that makes more sense now !! thanks a lot !

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

@amistre64 sry 1 more question consider a deck of 53 cards, which includes a joker (wild card). The possible poker hands are: (a) Five-of-a-kind (Four of one kind, plus a joker) (b) Straight Flush (ve cards same suit and in consecutive order). (c) Four-of-a-Kind (4,1) (d) Full House (3,2) (e) Flush (f) Straight (g) Three-of-a-Kind (3,1,1) (h) Two Pairs (2,2,1) (i) One Pair (2,1,1,1)

OpenStudy (anonymous):

(j) High Card: This occurs when none of the above hands is made. For example, a Four-of-a-Kind hand can be made in two ways: Without joker: four of one kind plus one of another kind. With a joker: three of one kind, plus a joder, plus one of another kind. Compute the number of ways to get each hand, and the probability. Check your work by adding the number of ways for all cases. The total must be C(53;5).